OpenStudy (anonymous):

Michael is three times older than his sister Angela. In 1 year, Michael will be twice as old as Angela. How old are the children today?

OpenStudy (ganpat):

Angela = x, Michael = 3x In 1 year, Michael = 2 (x +1) I can get this two equations only.. I guess, something is missing..

OpenStudy (anonymous):

oh okay, thank you though :)

OpenStudy (anonymous):

u want any more help now

OpenStudy (anonymous):

Well I wanted to finish the question cause the other guy only gave me one part I think

OpenStudy (anonymous):

do u want me to give it

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Angela = x, Michael = 3x In 1 year,Angela=(x+1) , Michael = 2 (x +1) 3x+1=2x+2 3x-2x=2-1 x=1 years and 3x=3 years

OpenStudy (anonymous):

medal please

OpenStudy (ganpat):

lol !!!!!

OpenStudy (ganpat):

nice attempt dude !!

OpenStudy (anonymous):

@brook_27 are u done now

OpenStudy (anonymous):

LOL thanks! :) wanna help me with one more? aha

OpenStudy (anonymous):

A chemist needs 10L of 21% salt solution. The chemist has two salt solutions availible at 15% and 25% salt. Write a solve a linear system to find the volume of each solution that needs to be combined to make the mixture

OpenStudy (anonymous):

@brook_27 anything for u and your homework

OpenStudy (anonymous):

Thank you soo much ! <3

OpenStudy (zarkon):

you should post new problems in a different thread

OpenStudy (anonymous):

@brook_27 i didnt get it

OpenStudy (anonymous):

Let M and A be the ages of Michael and Angela today. Solve the following for M and A:\[\{M=3A,M+1=2(A+1)\}\]\[\{M=3 \text{ years},A=1 \text{ year}\} \]

OpenStudy (ganpat):

You just need equation i guess, 15x/100 + 25 (10 - x)/100 = 10 (21)/100 15x + 25 (10 - x) = 10 (21)

OpenStudy (anonymous):

what are u doing @Ganpat

OpenStudy (ganpat):

got the linear equation for the salt solution.. x liter of 15 % + (10-x) liter of 25 % = to give 10 liters of 21 %

OpenStudy (anonymous):

I think I have to convert the percentage into decimals =l but thanks !

OpenStudy (anonymous):

@brook_27 just ask some other questions

OpenStudy (ganpat):

percentage is nothing but ratio to 100.. so 15 percent = 15/100

OpenStudy (anonymous):

okay. what about -> Flying into the wind, a plane take 6h to fly 3000km. On the return flight, with the same wind, the plane takes 5h to complete the trip. How fast does the plane fly without any wind, and how fast was the wind blowing?

OpenStudy (anonymous):

Let x and y be the volume of the 15% and 25% salt solutions respectively. Solve the following equations for x and y:\[\left\{x+y=10,\frac{15}{100}x+\frac{25}{100}y=\frac{21}{100}10\right\}\]\[\{x=4 L,y=6 L\} \]

OpenStudy (anonymous):

omg thanks !

OpenStudy (anonymous):

r * t = d Let s and w be the speed of the plane in still air and w be the wind speed. Solve the following for s and w:\[\{(s-w)6=3000,(s+w)5=3000\}\]\[\{s=550 \text{ km}/\text{hr},w=50 \text{ km}/\text{hr}\} \]